In a geometric sequence, a3=9, a6=9/8, find a 12
Answer by richwmiller(17219)

Tn = t * r^(n - 1)
9 = t* r^2
9/8 = t * r^5
t=36 r=1/2
T12 = 36 * 1/2^(12)
t12=9/1024